OpenStudy (anonymous):

Express in terms of logarithms x,y,z or w, (a) log ↓3 (xyz); (b) log ↓3 (xz/y); (c) log ↓4 ³√y

OpenStudy (mertsj):

The log of the product is equal to the sum of the logs.

OpenStudy (mertsj):

\[\log_{13}xyz=\log_{13}x+\log_{13}y+\log_{13}z \]

OpenStudy (mertsj):

The log of the quotient is equal to the difference of the logs.

OpenStudy (mertsj):

\[\log_{13}(\frac{xz}{y})=\log_{13}x+\log_{13}y-\log_{13z} \]

OpenStudy (mertsj):

\[\log_{14}y ^{\frac{1}{3}}=\frac{1}{3}\log_{13}y \]

OpenStudy (anonymous):

I'm sorry. I did not know how to denote Sub 3 using Alt Keys. It should read: log sub 3. (Sub 4 should also be sub 3) Hope this helps. We just started chapter on logarithms.

OpenStudy (anonymous):

BTW... thank you.